### v1

# Unfolding the kagome lattice to improve understanding of ARPES in CoSn

Abstract
Metallic kagome lattices are attracting significant attention as they provide a platform to explore the interplay between topology and magnetism. Angle-resolved photoemission spectroscopy (ARPES) plays a key role in unraveling their electronic structure. However, the analysis is often challenging due to the presence of multiple bands near the Fermi level. Indeed, each orbital generates three bands in a kagome lattice due to its three inequivalent sites, which soon becomes complicated if many orbitals are present. To address this complexity, using ARPES matrix elements can be highly beneficial. First, band symmetry can be determined through selection rules based on light polarization. We emphasize that, in kagome lattices, symmetry is not only determined by the orbital character but also by the relative phase between the three sublattices. Additionally, interference between the three sublattices leads to a strong modulation of ARPES intensity across neighboring Brillouin zones. We show how unfolded band calculations capture these modulations, helping with band identification. We apply these ideas to CoSn, whose simple structure retains the key features of a kagome lattice. Taking advantage of these two effects, we isolate the dispersion of each band and discuss novel correlation effects, selectively renormalizing the bands crossing the Fermi level and shifting the others.
I Introduction
Metallic compounds with kagome lattices offer an exceptional means to study the interplay between a topological band structure, inherent to the kagome lattice Guo and Franz (2009); Mazin et al. (2014), and magnetism, ubiquitous in 3d kagome systems Yin et al. (2022); Wang et al. (2024). Interesting properties of magnetic kagome materials include a large anomalous Hall effect in the magnetic Weyl semimetal Co3Sn2S2 Liu et al. (2018), a Chern quantum phase in TbMn6Sn6 Yin et al. (2020), and possibly large pseudomagnetic fields in strained FeSn Zhang et al. (2023). Other electronic instabilities, such as exotic charge density waves in the non-magnetic AV3Sb5 Guo et al. (2024) and magnetic FeGe Teng et al. (2022), are also under intense investigation. ARPES plays a crucial role in characterizing the electronic structure of these materials Ye et al. (2018); Kang et al. (2020a, b); Lin et al. (2020); Liu et al. (2020); Hu et al. (2023); Peng et al. (2024), allowing visualization of Dirac cones, van Hove singularities, and flat bands, characteristic of the kagome band structure. Moreover, it is well suited to characterize the strength of electronic correlations, appearing, for example, as renormalization of the calculated band dispersion. Most metallic kagomes are found to be weakly correlated, with renormalization factors less than 2, but some exhibit more correlations, like CsCr3Sb5 Peng et al. (2024). In these intermetallic kagome systems, a problem is often that the weak crystal field does not significantly lift the degeneracy of the five orbitals. This leads to a rather complex situation with 15 bands often strongly overlapping near the Fermi level (5 per orbital times the 3 kagome sites).
Fortunately, there are some ways to separate bands in an ARPES experiment. The most common method is to use selection rules associated with linear polarization to select orbitals that are either odd or even with respect to mirror planes containing high symmetry directions Damascelli et al. (2003); Moser (2023). In single-site systems, this depends solely on the orbital parity with respect to that mirror plane. In multi-site systems (e.g., non-Bravais lattices), such as the kagome lattice, this also depends on the relative phases of the orbitals at different sites Brouet et al. (2012), which is not always properly taken into account in the ARPES literature. We will detail these selection rules for a kagome lattice.
Another potentially useful selection rule is the intensity modulation of the bands in different Brillouin Zones (BZ), due to interference between the atoms of the motif. This effect is similar to the form factor in structural experiments, which causes extinctions of some Bragg peaks in non-Bravais lattices. It naturally appears in ARPES matrix elements Nishimoto et al. (1996); Moser (2023), but it is much less utilized. For a 2-site motif, there will typically be one bonding and one antibonding band. The interference is constructive only for the bonding band, yielding apparently different dispersion in the first and second BZ. This can, for example, be observed in perovskites with rotated octahedra, like iridates Martins et al. (2018) or ruthenates Sohn and Kim (2022), but also with some more subtleties in graphene Liu et al. (2011) or iron pnictides Brouet et al. (2012). The case of the 3 sites lattice has been much less studied, and it is less predictable. We show here that unfolded calculations Dirnberger et al. (2021); Rubel et al. (2014) can easily predict these effects in complex systems. The relevance of unfolded calculations to analyze ARPES was already indicated for supercell systems Ku et al. (2010). More surprisingly, it can also be applied to systems like graphene Lee et al. (2018), where the definition of supercell is less intuitive. We show that this can be rationalized within an extended zone scheme. We use here unfolded calculations to obtain the spectral weight of the different bands in different BZ. Although it is not the only contribution to ARPES matrix elements, it often has a significant impact.
We illustrate the benefit of taking these two points (polarization and the use of different BZ) into account with the example of CoSn, a relatively simple metallic and non-magnetic kagome system. Its electronic structure was already mapped out by ARPES Liu et al. (2020); Kang et al. (2020a); Cheng et al. (2023), but we pay particular attention to the role of matrix elements. By employing different polarizations along high-symmetry directions, we achieve a complete separation between odd and even bands. We further observe strong modulation of the bands’ intensities in different BZs, well predicted by unfolded calculations. We finally discuss the impact of correlations on each band and suggest different behaviors for the bands crossing the Fermi level and the filled ones at higher binding energies. This methodology should be increasingly helpful for more complex kagome systems that deviate more strongly from DFT calculations.
II Methods
Single crystals were grown from a Sn-flux method, as in reference Liu et al., 2020. Their structure was characterized by X-ray diffraction, their composition by Energy Dispersion X-ray analysis and their electronic properties by transport measurements. ARPES experiments were carried out on UHV-cleaved single crystals at the CASSIOPEE beamline of SOLEIL synchrotron and the BLOCH beamline of MAX-IV synchrotron, at low temperatures (T15 K) and with an overall resolution better than 15meV.
DFT calculation were done using the full-potential linear augmented plane wave method implemented in the WIEN2k software Blaha et al. (2001). The experimental crystal structure (space group P6/mmm, No. 191) of CoSn with lattice constants Å and Å was used in the calculations.
III Band parity in the kagome lattice
III.1 Tight-binding model
The three atoms triangular motif sketched in Fig. 1(a) is the hallmark of the kagome lattice. For each orbital, three bands will be created, in which the three atoms are combined with coefficients where j indicates the atomic site at position .
Let us label the central atom as 1 and the two orthogonal axes as and . The two planes and , which are perpendicular to the kagome plane, are mirror planes of the structure. They correspond respectively to K and M directions of the BZ (see inset in Fig. 1). One can see that reflection with respect to these mirror planes exchanges atoms 2 and 3 while leaving atom 1 unchanged. Therefore, the three bands created from a single orbital must combine atoms 2 and 3 in a symmetric () or antisymmetric () way to respect the mirror planes. In the first case, a contribution from atom 1 can be added (in-phase or out-of-phase), while in the second case, it must be zero.
We take as an example a orbital at site 1, which is even with respect to the mirror plane and odd with respect to the mirror plane. To preserve the symmetry of the kagome lattice, the orbital must rotate by 120∘ at sites 2 and 3. This is achieved by taking an appropriate combination of and orbitals Kim and Liu (2023). In other words, we consider the even component of the / combination. The typical dispersion of the three bands obtained from a tight-binding (TB) model with only first-neighbor couplings Guo and Franz (2009) is shown in Fig. 1(b). We neglect here spin-orbit coupling (SOC) that will open a gap where two bands touch at or K Guo and Franz (2009). This is crucial to analyse topological properties, but will not play a large role away from the crossing, where we will focus to identify the bands character. We further indicate on the figure the relative contributions of the three atoms in the first Brillouin zone, and we use color to represent the band parity with respect to mirror planes. The parity matches that of the orbital when sites 2 and 3 are in phase, and it is opposite when they are out-of-phase. Notably, forming a Dirac cone at K requires bands of opposite parity. However, it is often incorrectly assumed in ARPES studies of kagome compounds that band parity can be inferred directly from the dominant orbital character (leading to a wrong application of the selection rules that we describe below).
III.2 CoSn calculated band structure
CoSn has a relatively simple structure with kagome planes stacked vertically on top of each other and separated by two Sn2 buffer layers Sales et al. (2021). This preserves the characteristic features of the simple TB model of the kagome plane, especially Dirac cones and relatively flat bands, as can be seen in the DFT calculation Liu et al. (2020) at of Fig. 2(a). More precisely, each orbital displays a 3-band structure with very well-defined orbital character (see symbols). We label 15 the bands built from / orbitals and 710 bands from /, observable in the chosen energy window. The orbitals appear only below eV at =0 and will be ignored for the rest of the discussion (at =, a strong three-dimensional dispersion brings it much closer to the Fermi level). Two Dirac cones are formed around eV for / (bands 3 and 4) and / (bands 8 and 9). Two nearly flat bands are present near eV at for (band 7) and (band 2), with a dispersion of about 0.2 eV, much smaller than the 1 eV dispersion of the Dirac bands.
The three bands are clearly visible in this energy window, as highlighted by the marker size in Fig. 2(b). Some residual weight is observed along bands 1 and 2, which are predominantly from , evidencing hybridization between the two. The symmetry along M and K is described by the point group. This symmetry is characterized by two perpendicular mirror planes: the kagome plane itself and the perpendicular plane containing the considered symmetry direction. Since the symmetry with respect to this latter mirror plane defines the ARPES selection rules Damascelli et al. (2003), we group in Fig. 2(b) the four irreducible representations of into two groups: those that are even (red) and those that are odd (blue) with respect to this mirror plane. One can check that they follow the same rules as in the TB model of Fig. 1(b).
III.3 CoSn measured band structure with ARPES
We now turn to the comparison with actual ARPES experiments. In Fig. 3(a), we show a Fermi Surface for a CoSn sample lightly doped with Fe (5%), taken with linear horizontal (LH) polarization and at 126 eV photon energy, corresponding to =0 Kang et al. (2020a). In ARPES, polarization selection rules apply when the photoelectron momentum lies in high-symmetry planes, in this case and containing either K or M depending on the sample orientation. If the polarization is even (odd) with respect to this plane, only even (odd) bands will have non-zero matrix elements Damascelli et al. (2003); Moser (2023). Fig. 3(b) sketches the experimental geometry with the orientation of LH and LV polarization in our experiment. As LH lies within the plane, it is even with respect to it, and even bands will be detected along . As LH has both even and odd components with respect to , no straightforward selection rules can be applied. On the other hand, LV polarization is along and perpendicular to , selecting even bands along and odd ones along . Note that these selection rules only predict if a band can be observed and not if its intensity will be strong.
In Fig. 3(c-f), we show measurements along high symmetry directions, where strictly even and odd configurations are realized, either by switching polarization (Fig. 3(d) vs Fig. 3(c), along K) or the sample orientation (Fig. 3(f) vs Fig. 3(e) for M). We recall above each image the bands expected to be observed according to parities defined in Fig. 2(b). One can check that these selection rules are very well observed with indeed a perfect separation between even and odd bands, as indicated by numbers.
We also observe that the bands are almost systematically observed in only one of the two BZs. At first, this could appear puzzling, as it seemingly breaks the periodicity expected for the kagome lattice. This contrast between different BZs has often been noted in ARPES of kagome systems,but, to our knowledge, it has never been systematically explained.
IV Unfolding the kagome lattice
We show here that the expected interferences between the three kagome sites explain this behavior well. We first return to the TB model of the kagome lattice presented in Section II.A. In an extended zone scheme, we consider the wave vector , where is defined in BZ1 and is a reciprocal wave vector of the kagome lattice. From Eq. (1), we see that a phase factor [] will appear between values at and , which may be different from one in two neighboring BZ, revealing the hidden periodicity in the motif (see Appendices A and B). A triangular lattice of periodicity would define the larger BZ, represented as a dotted line in Fig. 4(a). In this BZ, there are two inequivalent sets of high-symmetry points, indicated as , K, M and ’, K’, M’ in the figure. Three inequivalent paths are obtained by shifting the path at [red, at ] by (green) or (blue). and are reciprocal wave vectors of the kagome BZ, but not of this larger BZ. In fact, atoms are dephased by along these paths, yielding for atoms 2 and 3 along and for atom 2 only along .
This will directly impact ARPES measurements in different BZs, as it was shown for other multisite systems like graphite or graphene Nishimoto et al. (1996); Moser (2017) that the intensity of the ARPES bands is modulated by the sum of the TB coefficients (see also the Appendix C).
Within the TB framework, this sum can easily be calculated and is represented by marker size in Fig. 4(b-d), for each inequivalent path. One can see that only one band is strong for each of the paths, specifically the one with the three atoms in phase (see also Appendix A, which describes the same effect as a function of the hidden periodicities). This emphasizes the need to measure ARPES at least along these three paths to observe all bands. In contrast, this is a means to better identify the nature of the band observed.
More generally, when a tight-binding (TB) model is either intractable or fails to capture the relevant physics, the electronic structure can be computed at the DFT level. The resulting wave functions can then be “unfolded” Dirnberger et al. (2021); Popescu_PRB_85_2012; Rubel et al. (2014) by projecting the states obtained in the true unit cell onto the reciprocal space of an alternative unit cell that accounts for additional periodicities
Such procedures are now commonly used to compare theoretical calculations with ARPES measurements when a supercell arises from an electronic ordering, such as antiferromagnetism or charge density waves. The resulting folded bands, with periodicity imposed by electronic order, typically exhibit weak intensity, in fact proportional to the strength of the perturbation Voit et al. (2000); Brouet et al. (2004). This approach can also be beneficial in calculations that involve large unit cells, for example, to simulate doping effects Berlijn et al. (2012).
We propose here to use it to analyze the kagome lattice, where the origin of the subperiodicities is intrinsic to the non-Bravais lattice type (see Appendix A). To unfold the band structure in the large BZ (dotted line in Fig. 4(a)), we use the Fold2Bloch module Rubel et al. (2014, 2023) and a transformation matrix . Figure 5 displays this calculation unfolded along the three different paths identified in Fig. 4(a). The size of the marker gives the unfolded weight corresponding to Eq. 2 and the colors refer to the orbital character, as in Fig. 2. There are well defined differences expected for spectral weights in different BZ, which will be the largest when all atoms are in-phase. For example band 1 is very strong in (a), while bands 3 and 4 are nearly zero. The opposite is true in panel (b) and (c), but differently for M and K. Generally, the unfolding effect is less clear on / orbitals, because they are more strongly hybridized with each other, blurring the differences in atomic composition ( and having opposite parities, they can only be hybridized for different atomic combinations).
The data in Fig. 3(c-f) should be compared to Fig. 5(a) for BZ1 and Fig. 5(b) for BZ2. Already in the FS, it is clear that the intensity is much larger in BZ1 than in the neighboring BZ2, evidencing these modulations. The circular FS is formed by band 1, the only band crossing at =0. Its larger intensity in BZ1 corresponds very well to prediction of the unfolded calculation. In all dispersion images, there is almost no trace of bands 3 and 8 at 1, as also predicted by unfolding. On the other hand, bands 3 and 8 appear strongly at 2 along K [Fig. 3(c-d)] and more weakly along M [Fig. 3(e-f)] in good agreement with Fig. 5(b). Both bands 3 (/) and 8 (/) form Dirac cones at K for similar energies (black arrow), so that separating them as in Fig. 3(c-d) is useful to define their respective dispersions and the energies of the Dirac crossings. On the contrary, band 7, having a flat part near EF at K, is present in Fig. 3(c) in both BZ, although with stronger weight on slightly different windows. This also corresponds well to the smaller unfolding effect we noted for /. The fact that its intensity is distributed over all zones may explain its relatively weaker intensity.
This shows that many strong variations in band intensity across neighboring Brillouin zones (BZs) can be explained by unfolding effects. However, a few deviations are noted compared to the calculated spectral weights. Band 4 appears significantly more intense in BZ1 than in BZ2 [this is particularly evident along the M path in Fig. 3(f)], whereas the opposite trend is predicted in Fig. 5. Band 2 fades away along KM’ in Fig. 3(b), as expected, but becomes very strong along M’K’, whereas a strong weight is instead expected along MK in Fig. 5. These differences may reflect the role of other matrix element effects, which can mask the variation arising from sublattice interference. On one hand, the polarization cross section can vary over -space and, on the other hand, the photon energy cross section can be quite different for different bands. For example, in CoSn, the intensity of band 7 is weak at 126 eV but much stronger at 70 eV (not shown), while the opposite is true for band 2. More deeply, the ARPES intensity is proportional to the unfolding weight of Eq. 3 assuming common approximations (see Appendix C) that can also have their limit.
V Electronic correlations in CoSn
Figure 6(a) gives an overview of almost all the bands, by adding images measured in different experimental conditions in pure CoSn. Five bands are close together at around 0.4 eV and it is useful to use the combination of unfolding and parity outlined in Fig. 3 to separate them. Although SOC mixes the bands in the crossing region, blurring there the definition of parity and orbital character, they remain well enough defined away from to perform this analysis. On the other hand, this mixing is at the origin of the non-trivial topology of CoSn, which was precisely studied in ref. Kang et al., 2025. A zoom on the 0.6 eV region below is shown in Fig. 6(b,c) with models of the experimental dispersion as symbols. Our findings are in very good agreement with previous ARPES studies of CoSn Liu et al. (2020); Kang et al. (2020a); Cheng et al. (2023), which all find a bottom position of band 1 near 0.2 eV, the flat band 7 near 0.05 eV at K and the Dirac cones around 0.8 eV. Qualitatively, this agrees well with the calculated band structure, including SOC, which is shown as thin lines in Fig. 6(a). Quantitatively, the calculated dispersions are in fact shifted up by 0.18 eV, evidencing some deviation. The calculated bottom of band 1 at 0.37 eV is significantly different from the experimental value.
In principle, one would expect electronic correlations to induce a renormalization of the dispersions rather than an energy shift, as the latter does not conserve the number of carriers. Indeed, the Fermi level crossing of band 1 is not well described in Fig. 6(a) (the green arrows indicate the deviation). In Fig. 6(b), band 1 is renormalized by a factor of 2, which describes very well both the bottom of the band and its . This indicates that our data agree with the expected CoSn stoichiometry at . However, if all bands are renormalized by a factor of 2, as done in Fig. 6(b), the agreement worsens at higher binding energies. Band 3 is 0.05 eV lower in experiment than in this fit, and it is more dispersive [see green arrow in Fig. 6(b)]. At even higher binding energies [see Fig. 6(a)], band 4 clearly disperses from 1.35 eV at to 0.8 eV at K, which cannot be described by the renormalized band structure. We observe that the splitting between band 1 and 3 is 0.1 eV, exactly as in the unrenormalized band dispersion with SOC. This suggests that only the band crossing the Fermi level is renormalized. As it is tied by symmetry to other bands (band 1 and band 2 are only separated by the SOC gap, for example), renormalizing band 1 necessarily raises band 2 and the bands strongly hybridized with it. In Fig. 6(c), we renormalize band 1 by 2 and shift up all other bands by 0.18 eV, which gives a reasonable fit, although further orbital-dependent adjustments may be necessary. We note that a rather large range of renormalization values was given for CoSn, between 1.2 Kang et al. (2020a); Cheng et al. (2023) and 2 Liu et al. (2020), probably depending on which features were predominantly fitted.
VI Conclusion
We have shown how bands can be separated according to their parity through polarization analysis adapted to the kagome lattice and how systematic variation of the ARPES band intensities in neighboring Brillouin zones can be captured by relatively simple unfolded DFT calculations. This provides additional tools to characterize the nature of a band beyond its dispersion. In the non-magnetic kagome compound CoSn, isolating each dispersion suggests that the bands are not identically renormalized, as usually assumed, but that the band crossing the Fermi level is renormalized by a factor of 2, pulling up all other bands by as much as 0.2 eV. This manifestation of electronic correlations appears quite different from what was observed in other multi-orbital systems, such as iron pnictides Yi et al. (2017), where the renormalization rather seems to depend on orbital character. This could be interesting to investigate further.
Following this idea, if many bands cross the Fermi level with different occupied bandwidths, one could expect a rather complex pattern of shifts compared to DFT calculations. In this case, it becomes crucial to have additional ways to identify the nature of a band to make sense of a comparison between ARPES and DFT. This is probably what happens in the magnetic FeSn, where many bands cross the Fermi level and the comparison with DFT calculations is indeed not straightforward at all anymore Ghimire and Mazin (2020). Symmetry-resolved unfolded calculations should help achieve a rational understanding of the ARPES measurement.
Appendix A From triangular to kagome lattice
To unfold the kagome lattice, we used the intermediate step of a fictitious triangular lattice. We detail here its relationship with the kagome lattice. The dispersion in a tight-binding model for a 2D triangular lattice of side a with nearest-neighbor hoppings has the following form.
In Fig. 7, we display this dispersion in the kagome unit cell (red line), as well as its translated by =(0,2/a) (green) and =(/a,/a) (blue), as defined in Fig. 4. The resemblance with the kagome dispersion is already apparent. In particular, it is clear that the strongest weight along the three paths indicated in Fig.4 follows quite closely the band of the triangular lattice translated by the corresponding wave vector. Let us note that the translation , also included in an unfolded calculation, is equivalent by symmetry to .
To obtain the kagome lattice, we "remove" progressively the site at the center of the hexagon by tuning the hopping with this site to zero. We obtain the kagome dispersion shown as dotted black lines, where gap have appeared, where some of the translated bands touched. This is also the regions where the unfolded weight is most distributed between the bands.
Appendix B Extended zone scheme
Standard band unfolding techniques rely on a known relationship between the supercell and an assumed primitive cell Wang et al. (1998); Boykin and Klimeck (2005). Identifying the parent cell is often straightforward in systems with structural distortions (e.g., octahedral tilting Rubel et al. (2023)), magnetic order, or disordered alloys modeled via random atomic substitutions in a supercell Rubel et al. (2014). However, in some cases, superlattice effects are more subtle and difficult to disentangle. A prominent example is quasicrystals, which display sharp Bragg peaks despite lacking conventional long-range periodicity Bindi et al. (2009). The extended zone scheme provides a way to represent the electronic structure without assuming a specific primitive cell or Brillouin zone. This approach has previously been applied to aperiodic systems to extract a quasi-dispersion relation Niizeki and Akamatsu (1990); Mayo et al. (2020).
Let us assume that we wish to construct an extended zone representation for an eigenstate of an aperiodic structure expressed in a tight-binding basis
Here, labels the atomic site positions, is the total number of sites, denotes the site amplitude, and is an atom-centered basis function. Due to the lack of translational symmetry, this state does not possess a well-defined crystal momentum. The spectral function in the extended zone scheme is obtained via the generalized structure factor Niizeki and Akamatsu (1990)
This result is analogous to the diffraction pattern of a quasicrystal, where serves as a site-dependent scattering amplitude Niizeki and Akamatsu (1990).
In periodic solids, an electronic eigenstate with wave vector can be expanded in plane waves
where is the unit cell volume, and are the plane wave expansion coefficients. The summation is carried out over reciprocal vectors
with denoting the reciprocal lattice basis vectors. By analogy with Eq. (6), the spectral function can be written as
The Dirac delta function effectively selects a single plane wave component, satisfying , that contributes to the spectral weight. Alternative formulations of the spectral function in the extended zone, aimed at analyzing ARPES measurements, have also been proposed in the literature Kosugi et al. (2017); Mayo et al. (2020).
In ARPES, the measured intensity reflects the momentum-resolved spectral function near the surface region. To simulate this, we compute the spectral function projected onto a plane parallel to the surface (001)
The incoherent sum over plane wave coefficients in this expression implies that Fourier components with nonzero out-of-plane wave vectors are folded back into the projection plane. The Fermi surface within the plane in the extended zone representation can then be constructed as
Figure 8(a) shows a slice through the Fermi surface of CoSn calculated using the Vienna ab initio simulation package (VASP) Kresse and Hafner (1993); Kresse and Furthmüller (1996, 1996) and presented in the extended zone scheme. A circular pattern appears in the first Brillouin zone but is absent in the adjacent periodic image, reappearing only in the third shell. This doubling of periodicity in reciprocal space suggests a Brillouin zone that is twice as large, which corresponds to a perceived primitive cell that is twice as small in real space, as illustrated in Fig. 8(b). This result supports the validity of the unfolding method and confirms the appropriateness of the chosen primitive cell.
Appendix C ARPES matrix elements
Comprehensive treatments of ARPES principles can be found in textbooks Hüfner (2013); Tejeda and Malterre (2024) and papers Damascelli et al. (2003); Moser (2017); Yen et al. (2024). Applying Fermi’s golden rule, the photoemission intensity will depend on matrix elements of the form:
where describes the interaction between electron and photon, and are initial and final states. These are usually -particle states, but for our discussion, only the one-electron matrix element will matter. is usually developed as a perturbation in the dipolar approximation , although this is not sufficient for some cases (especially when the Hamiltonian is non-local Hwang et al. (2011)). The final state is usually taken as a plane wave of wave-vector , assuming the photoelectron is free to travel to the detector. Corrections to this approximation are often important to properly describe the dependence on photon energy or polarization Kern et al. (2023). Unfortunately, this often requires sophisticated calculations, which miss a simple connection with the properties of the initial state.
We consider the matrix element =. Using the TB expression given in Eq. 1 as initial state, we get.
To calculate the ARPES intensity, we follow Nishimoto et al., who derived Eq. (2) in the case of graphite Nishimoto et al. (1996). A similar derivation can be found elsewhere Moser (2017); Yen et al. (2024). The term in brackets essentially depends on the shape of the orbital at site . They are all the same in our case and we call =, where is a lattice site. Recalling the final state is a Bloch function of wave vector , we have .
If we rather write the initial state as a function of the sublattice periodicities (see Appendix A), the delta function selects a value =, leading to Eq. (3).